OpenStudy (anonymous):

A baker has a bag of flour that is 11% whole wheat and a bag of flour that is 63% whole wheat. How many cups of each type does the baker need to make 12 cups of a flour mixture that is 50% whole wheat? (1 point) 3 cups of the 11% flour and 9 cups of the 63% flour 9 cups of the 11% flour and 3 cups of the 63% flour 3 cups of the 11% flour and 3 cups of the 63% flour 9 cups of the 11% flour and 9 cups of the 63% flour

OpenStudy (anonymous):

hello any one there

OpenStudy (whpalmer4):

Think of this in terms of how much whole wheat there is in each cup. 1 cup of 11% whole wheat = 0.11 C whole wheat and 0.89 C other. 1 cup of 63% whole wheat = 0.63 C whole wheat and 0.37 C other. Let's say A represents the amount of the 11% ww flour you use, and B represents the 63% ww flour. Then we can write A + B = 12, because we want to make 12 cups of flour. We also know that we need to end up with 12 cups * 50% = 6 cups of whole wheat in those 12 cups, so we can write another equation to show how we get that.

OpenStudy (whpalmer4):

\(0.11 * A + 0.63 * B = 6\) Now we know two equations in two unknowns, and we can solve them by substitution or elimination. Do you know how to do that?

OpenStudy (anonymous):

yeah i think

OpenStudy (whpalmer4):

Alternatively, because we have our answer choices in front of us, you could just try plugging them into the second formula and see if they produce a correct result (both sides are equal). But you should really know how to solve them, because most of the time you won't have that option in real life :-)

OpenStudy (whpalmer4):

Try it, I'll check your answer.

OpenStudy (anonymous):

X=6/0.11A+0.63B

OpenStudy (whpalmer4):

Here are our two equations: \[A + B = 12\]\[0.11A + 0.63B = 6\] I would solve the first equation for A, and substitute that into the second equation: \[A = 12-B\]\[0.11(12-B)+0.63B = 6\]At this point, multiply both sides by 100 to get rid of the decimals: \[11(12-B) + 63B = 600\]expand the left side: \[132 - 11B + 63B = 600\]Solve for \(B\), then use \(A = 12-B\) to find \(A\).

OpenStudy (anonymous):

all i get is A=12-B

OpenStudy (whpalmer4):

\[132- 11B + 63B= 600\]Can you move the 132 to the right hand side?

OpenStudy (anonymous):

it will be B=9

OpenStudy (whpalmer4):

Yes!

OpenStudy (anonymous):

so th answer will be b or d

OpenStudy (anonymous):

9 cups of the 11% flour and 9 cups of the 63% flour

OpenStudy (whpalmer4):

Only one matches our problem setup. Look at what I wrote above about what A and B represented.

OpenStudy (whpalmer4):

Also, we only are making 12 cups of flour, so d) can't be correct, because that would make 18 cups.

OpenStudy (whpalmer4):

Similarly, c) can't be correct, because that only makes 6 cups of flour. I would take another look at choices a and b...

OpenStudy (anonymous):

so it will be 9 cups o the 11% flour and 3 cups of the 63% flour

OpenStudy (whpalmer4):

And look at our equation: even if we don't remember which type of flour is going to be 9 cups, we know that the percentages multiplied by 3 cups and 9 cups must add to 6 cups of whole wheat flour (the thing of interest in the mixture).

OpenStudy (whpalmer4):

Come on, read what I wrote up there again. Here is the relevant sentence: "Let's say A represents the amount of the 11% ww flour you use, and B represents the 63% ww flour. "

OpenStudy (whpalmer4):

if we have 9 cups of 11% flour, that means we have 9 * 0.11 = 0.99 cups of ww flour from that part, and 3*0.63 = 1.89 cups of ww flour from the other part. Does that add up to 6 cups of ww flour in the mixture?

OpenStudy (whpalmer4):

If we have 3 cups of 11% flour, that means we have 3*0.11 = 0.33 cups of ww flour from that part, and 9*0.63 = 5.67 cups of ww flour from the other part. Does that add up to 6 cups of ww flour in the mixture?

OpenStudy (anonymous):

already got the answer man